# Automatic Debiased Machine Learning for Covariate Shifts

### v4

@@ 2.1 Parameters of interest and debiased estimating equations
[…]
@@ 2.2 Debiased estimation with cross-fitting
$\alpha\mathopen{}\mathclose{{}\left(x}\right)=b\mathopen{}\mathclose{{}\left(x}\right)^{{\mathrm{\scriptscriptstyle T}}}\rho;\quad\operatorname{pen}(a)=r_{a}\sum_{j=1}^{p}\mathopen{}\mathclose{{}\left|\rho_{j}}\right|,$
@@
For each fold $\ell\,$of the cross-fitting, a debiased machine learner can be constructed as the sum of a plug-in estimator and a bias correction
@@ 2.3 Large sample theory with cross-fitting
a) The defining formula $m(Z,\gamma)$ is linear in $\gamma$. b) The training data $(Y_{t},X_{t})$ for $(t=1,...,T)$, and the target data $(Z_{i})$ for $(i=1,...,N)$, are each generated as independent and identically distributed random vectors, with the two samples mutually independent. c) The ratio of target to training data sizes, $N/T$ converges to $0\leq\xi<\infty$. d) The functional $m(Z,\gamma)$ is mean square continuous in $\gamma$, namely $E\{m(Z,\gamma)^{2}\}\leq C\mathopen{}\mathclose{{}\left\|\gamma}\right\|_{2}^{2}$ for some $C<\infty$. e) The representer $\alpha_{0}(X)$ and conditional variance $\text{var}(Y|X)$ are bounded.
@@ 2.4 Debiased estimation without cross-fitting via lasso
$\hat{V}=\frac{1}{N}\sum_{i=1}^{N}\{m(Z_{i},\hat{\gamma})-\bar{m}\}^{2}+\frac{N}{T}\frac{1}{T}\sum_{t=1}^{T}\tilde{\alpha}(X_{t})^{2}\{Y_{t}-\hat{\gamma}(X_{t})\}^{2},\quad\bar{m}=\frac{1}{N}\sum_{i=1}^{N}m(Z_{i},\hat{\gamma}),$
@@ 2.5 Large sample theory without cross-fitting
This condition simplifies the analysis considerably. Let $J$ denote a subvector of $(1,...,p)$, $\mathopen{}\mathclose{{}\left|J}\right|$ denote the number of elements of $J,$ $\beta_{J}$ be the vector consisting of $\beta_{Jj}=\beta_{j}$ for $j\in J$ and $\beta_{Jj}=0$ otherwise, and $\beta_{J^{c}}$ be the corresponding vector for $J^{c}$. Hence $\beta=\beta_{J}+\beta_{J^{c}}$.
@@
Sparse coefficients have a minimal eigenvalue: there is $c>0$, such that given any degree of sparsity $s=O\{n/\ln(p)\}$, with probability $1-o(1)$, the sparse restricted eigenvalue is bounded: $\min_{\mathopen{}\mathclose{{}\left|J}\right|\leq s}\min_{\mathopen{}\mathclose{{}\left\|\beta_{J^{c}}}\right\|_{1}\leq 3\mathopen{}\mathclose{{}\left\|\beta_{J}}\right\|_{1}}{\beta^{{\mathrm{\scriptscriptstyle T}}}\hat{Q}\beta}/{\beta_{J}^{{\mathrm{\scriptscriptstyle T}}}\beta_{J}}\geq c.$
@@ 3 Simulaton experiments
This polynomial is of order $Q=3$ over $K=6$-dimensional vectors $x$, includes cross terms, and is parameterized by $\delta$ and $\iota$. The outcome is generated as $Y_{t}=\gamma_{0}(X_{t})+\epsilon_{t},$ where $\epsilon_{t}\sim N(0,\sigma^{2})$ with $\sigma=0.1$. The components of $X_{t}$ are independently generated from a mixture of $N(0,1)$ and $U(-5,5)$ distributions with mixing probabilities $0.99$ and $0.01$, respectively. The shifted covariates $Z$ are generated from $N(\mu,1)$ with a mean shift $\mu=1.1\sigma$. The $\delta_{q}$ parameters are set as $\delta_{1}=1$, $\delta_{2}=0.7$, and $\delta_{3}=0.5$.
@@ 4 Difference-in-differences estimation of minimum wage effect on teen employment
We apply our methodology to estimate the average treatment effect on the treated using unconfounded difference-in-differences identification. We have data $(Y_{0,i},Y_{1,i},D_{i},X_{i})$ for $(i=1,...,n)$, which are independent and identically distributed copies of the random variables $(Y_{0},Y_{1},D,X)$. Here, $Y_{0}$ and $Y_{1}$ are outcomes in two time periods for the same unit, $D$ indicates whether the unit was treated at the end of the first period, and $X$ are control variables. Under the conditional parallel trends assumption (e.g., Roth et al., (2023)), the effect is identified as
$\theta_{0}=E\{\Delta Y-\gamma_{0}(X)\mid D=1\},\quad\gamma_{0}(X)=E\{\Delta Y\mid D=0,X\},\quad\Delta Y=Y_{1}-Y_{0}.$
@@ A.1 Asymptotic linearity
    $E\{(R_{1}^{\prime})^{2}\mid\hat{\gamma}\}$ $=E([E_{N}\{m(Z,\hat{\gamma}-\gamma_{0})\}-\int m(z,\hat{\gamma}-\gamma_{0})f(z)\mathrm{d}z]^{2}\mid\hat{\gamma})$
    $=\frac{1}{N}E(E_{N}[\{m(Z,\hat{\gamma}-\gamma_{0})-\int m(z,\hat{\gamma}-\gamma_{0})f(z)\mathrm{d}z\}^{2}]\mid\hat{\gamma})$
@@
    $=\frac{1}{T}E([\alpha_{0}(X)\{\gamma_{0}(X)-\hat{\gamma}(X)\}-\int\alpha_{0}(x)\{\gamma_{0}(x)-\hat{\gamma}(x)\}f(x)\mathrm{d}x]^{2}\mid\hat{\gamma})$
@@
    $E(R_{2}^{2}\mid\hat{\alpha})$ $=E\{(E_{T}[\{\hat{\alpha}(X)-\alpha_{0}(X)\}\{Y-\gamma_{0}(X)\}])^{2}\mid\hat{\alpha}\}$
    $=\frac{1}{T}E(E_{T}[\{\hat{\alpha}(X)-\alpha_{0}(X)\}^{2}\{Y-\gamma_{0}(X)\}^{2}]\mid\hat{\alpha})$
@@
    $E(|R_{3}|\mid\hat{\gamma},\hat{\alpha})$ $=E(|E_{T}[\{\hat{\alpha}(X)-\alpha_{0}(X)\}\{\gamma_{0}(X)-\hat{\gamma}(X)\}]|\mid\hat{\gamma},\hat{\alpha})$
    $\leq E(E_{T}[|\{\hat{\alpha}(X)-\alpha_{0}(X)\}\{\gamma_{0}(X)-\hat{\gamma}(X)\}|]\mid\hat{\gamma},\hat{\alpha})$
@@ A.2 Trimming
The following facts hold under the last hypotheses of Theorem 1 and 2, where $1_{(\cdot)}$ is one if ()($\cdot$)( ⋅ ) is true and zero otherwise:
@@
    $E(|E_{T}[\{\hat{\alpha}(X)-\alpha_{0}(X)\}^{2}]|\mid\hat{\alpha})$ $=E(E_{T}[\{\hat{\alpha}(X)-\alpha_{0}(X)\}^{2}]\mid\hat{\alpha})$
@@ A.3 Main result
    $=m(Z,\hat{\gamma}-\gamma_{0})+E\{m(Z,\gamma_{0})\}-E_{N}\{m(Z,\gamma_{0})\}+E_{N}\{m(Z,\gamma_{0}-\hat{\gamma})\}.$
@@ Appendix B Proof of Theorem 2
Recall that $(\hat{\beta},\hat{\rho})$ are the coefficients of $(\hat{\gamma},\hat{\alpha})$. Let $(\beta_{N},\rho_{N})$ be the coefficients of $(\gamma_{N},\alpha_{N})$, and define $\theta_{N}=\beta_{N}^{\top}Q\rho_{N}$. Let $\hat{\mu}=E_{T}\{Yb(X)\}$, $\hat{U}=\hat{\mu}-\hat{Q}\beta_{N}$, and $\hat{R}=\hat{M}-\hat{Q}\rho_{N}$.

### v5

@@ 2.1 Parameters of interest and debiased estimating equations
$\theta_{0}=\int E\{\gamma_{0}(d+1,Z)-\gamma_{0}(d,Z)\}\,d\mu(d);\quad\theta_{0}=\int E\{\partial_{D}\gamma_{0}(d,Z)\}\,d\mu(d),$
averaged over values of $d$ according to a prespecified measure $\mu$. These parameters identify the average incremental treatment effect on the potential outcome $Y(d)$ for a target population with covariates $Z$. Here, $m(z,\gamma)=\int\{\gamma(d+1,z)-\gamma(d,z)\}\,d\mu(d)$ and $m(z,\gamma)=\int\{\partial_{D}\gamma(d,z)\}\,d\mu(d)$. For other examples that fall within this framework, see, e.g., Chernozhukov et al., 2022a .
@@ 2.2 Debiased estimation with cross-fitting
$\alpha\mathopen{}\mathclose{{\left(x}}\right)=b\mathopen{}\mathclose{{\left(x}}\right)^{{\mathrm{\scriptscriptstyle T}}}\rho;\qquad\operatorname{pen}(a)=r_{a}\sum_{j=1}^{p}\mathopen{}\mathclose{{\left|\rho_{j}}}\right|,$
@@
[…]
@@
$\text{ }\hat{s}_{\ell\alpha}^{2}$ $=\frac{1}{T_{\ell}}\sum_{t\in I_{\ell}}\tilde{\alpha}_{\ell}(X_{t})^{2}\{Y_{t}-\hat{\gamma}_{\ell}(X_{t})\}^{2}.$
@@ 2.3 Large sample theory with cross-fitting
a) The defining formula $m(Z,\gamma)$ is linear in $\gamma$. b) The training data $(Y_{t},X_{t})$ for (t=1,,T)(t=1,...,T), and the target data $(Z_{i})$ for (i=1,,N)(i=1,...,N), are each generated as independent and identically distributed random vectors, with the two samples mutually independent. c) The ratio of target to training data sizes, $N/T$ converges to $0\leq\xi<\infty$. d) The functional $m(Z,\gamma)$ is mean square continuous in $\gamma$, namely $E\{m(Z,\gamma)^{2}\}\leq C\mathopen{}\mathclose{{\left\|\gamma}}\right\|_{2}^{2}$ for some $C<\infty$. e) The representer $\alpha_{0}(X)$ and conditional variance $\text{var}(Y|X)$ are bounded.
@@ 2.4 Debiased estimation without cross-fitting via lasso
[…]
@@ 2.5 Large sample theory without cross-fitting
This condition simplifies the analysis considerably. Let $J$ denote a subvector of (1,,p)(1,...,p), $\mathopen{}\mathclose{{\left|J}}\right|$ denote the number of elements of $J,$ $\beta_{J}$ be the vector consisting of $\beta_{Jj}=\beta_{j}$ for $j\in J$ and $\beta_{Jj}=0$ otherwise, and $\beta_{J^{c}}$ be the corresponding vector for $J^{c}$. Hence $\beta=\beta_{J}+\beta_{J^{c}}$.
@@
Sparse coefficients have a minimal eigenvalue: there is $c>0$, such that given any degree of sparsity $s=O\{n/\ln(p)\}$, with probability 1o(1)1-o(1), the sparse restricted eigenvalue is bounded: $\min_{\mathopen{}\mathclose{{\left|J}}\right|\leq s}\min_{\mathopen{}\mathclose{{\left\|\beta_{J^{c}}}}\right\|_{1}\leq 3\mathopen{}\mathclose{{\left\|\beta_{J}}}\right\|_{1}}{\beta^{{\mathrm{\scriptscriptstyle T}}}\hat{Q}\beta}/{\beta_{J}^{{\mathrm{\scriptscriptstyle T}}}\beta_{J}}\geq c.$
@@ 3 Simulaton experiments
This polynomial is of order $Q=3$ over $K=6$-dimensional vectors $x$, includes cross terms, and is parameterized by $\delta$ and $\iota$. The outcome is generated as $Y_{t}=\gamma_{0}(X_{t})+\epsilon_{t},$ where $\epsilon_{t}\sim N(0,\sigma^{2})$ with $\sigma=0.1$. The components of $X_{t}$ are independently generated from a mixture of $N(0,1)$ and U(5,5)U(-5,5) distributions with mixing probabilities $0.99$ and $0.01$, respectively. The shifted covariates $Z$ are generated from $N(\mu,1)$ with a mean shift $\mu=1.1\sigma$. The $\delta_{q}$ parameters are set as $\delta_{1}=1$, $\delta_{2}=0.7$, and $\delta_{3}=0.5$.
@@ 4 Difference-in-differences estimation of minimum wage effect on teen employment
We apply our methodology to estimate the average treatment effect on the treated using unconfounded difference-in-differences identification. We have data $(Y_{0,i},Y_{1,i},D_{i},X_{i})$ for (i=1,,n)(i=1,...,n), which are independent and identically distributed copies of the random variables $(Y_{0},Y_{1},D,X)$. Here, $Y_{0}$ and $Y_{1}$ are outcomes in two time periods for the same unit, $D$ indicates whether the unit was treated at the end of the first period, and $X$ are control variables. Under the conditional parallel trends assumption (e.g., Roth et al., (2023)), the effect is identified as
$\theta_{0}=E\{\Delta Y-\gamma_{0}(X)\mid D=1\},\qquad\gamma_{0}(X)=E\{\Delta Y\mid D=0,X\},\qquad\Delta Y=Y_{1}-Y_{0}.$
@@ A.1 Asymptotic linearity
[…]
@@
[…]
@@
[…]
@@
[…]
@@ A.2 Trimming
The following facts hold under the last hypotheses of Theorem 1 and 2, where $1_{(\cdot)}$ is one if $(\cdot)$ is true and zero otherwise:
@@
[…]
@@ A.3 Main result
[…]
@@ Appendix B Proof of Theorem 2
[…]